# Salient Region-Guided Spacecraft Image Arbitrary-Scale Super-Resolution Network

### v1

@@ IV-B Comparison Results on Spacecraft ISAR Image Dataset
TABLE II: Quantitative results on spacecraft optical image dataset. The best and the second best results are marked in bold and underline.
Methods $\times$4 $\times$6 $\times$8 $\times$12
@@
SSRN [12] 33.11 0.6967 - - 31.99 0.6779 - -
SSNet [23] 32.56 0.6607 - - 31.30 0.6494 - -
VapSR [31] 32.43 0.6876 - - 30.64 0.6501 - -
LKFN [33] 32.96 0.7011 - - 31.73 0.6813 - -
ELAN [35] 32.83 0.6948 - - 31.88 0.6705 - -
SRConvNet [37] 32.95 0.6908 - - 31.56 0.6594 - -
LIIF [14] 29.03 0.6425 32.63 0.6772 29.24 0.6407 26.78 0.6313
LTE [15] 32.14 0.6309 32.70 0.6762 31.03 0.5895 30.19 0.5537
LM-LTE [43] 32.17 0.6314 32.68 0.6763 31.07 0.5896 30.23 0.5530
SGSASR(Our approach) 33.57 0.7132 32.72 0.6851 32.09 0.6708 31.17 0.6531
Figure 7: Visual comparison for fixed-scale SR methods on spacecraft optical image dataset. (a): ×$\times 4$. (b): ×$\times 8$.
Figure 8: Visual comparison for arbitrary-scale SR methods on spacecraft optical image dataset. (a): ×$\times 4$. (b): ×$\times 6$. (c): ×$\times 8$. (d): ×$\times 12$.
IV-C Comparison Results on Spacecraft Optical Image Dataset
In this section, we compare SGSASR with the nine methods mentioned in Section IV-B on spacecraft optical image dataset. All methods are retrained on spacecraft optical image dataset for fair comparison. The quantitative results are presented in Table II. We can observe that SGSASR still outperforms fixed-scale SR methods in terms of PSNR performance on spacecraft optical image dataset. Specifically, compared with SSRN, SGSASR respectively achieves improvements of 0.46 dB and 0.10 dB at two SR scales ($\times$4, $\times$8). We can also observe that SGSASR significantly achieves superior PSNR performance at various SR scales compared with other arbitrary-scale SR methods on spacecraft optical image dataset. Specifically, compared with LM-LTE, SGSASR respectively achieves improvements of 1.40 dB, 0.04 dB, 1.02 dB and 0.94 dB at four SR scales ($\times$4, $\times$6, $\times$8, $\times$12).
In addition, we present visual comparison results for fixed-scale SR methods in Fig. 7 and for arbitrary-scale SR methods in Fig. 8. Overall, compared with other fixed-scale SR methods, SGSASR achieves comparable or even better restoration of spacecraft structural details on spacecraft optical image dataset. In contrast to other arbitrary-scale SR methods, the spacecraft optical SR images generated by SGSASR not only more effectively restore the structure of the spacecraft, but also avoid introducing noise into the black background regions, which clearly demonstrates the superiority of SGSASR.
TABLE III: Effect of Proposed Modules in SFEEM
@@ IV-C Comparison Results on Spacecraft Optical Image Dataset
TABLE IV: Effect of Feature Aggregation Methods
@@
TABLE V: Effect of Lantent Modulation in LMASSR
@@
IV-D Ablation Studies
IV-D1 The Effect of Proposed Modules in SFEEM
@@
To verify the effectiveness of the two main contributions proposed in this paper—–SCRRB and AFFEM, we conduct the incremental ablation experiments on the spacecraft ISAR image dataset. The results are shown in Table III. The first row describes the Baseline model, which directly uses NAFNet [47] as the spacecraft image feature encoder. In the second row, we add the SCRRB, which yields a performance gain of 0.27 dB, since the SCRRB can better identify the core salient region in spacecraft images. In the third row, we further add the AFFEM, which yields an additional performance gain of 0.08 dB, since AFFEM can aggregate the spacecraft region features with general image features more flexibly.
IV-D2 The Effect of Feature Aggregation Methods
@@
To verify the advantage of our AFFEM, we compare it with other feature aggregation methods such as summation and concatenation. The results are shown in Table IV. We can observe that AFFEM achieves the best performance, demonstrating its effectiveness in adaptively aggregating the spacecraft region features with general image features.
IV-D3 The Effect of the Lantent Modulation in LMASSR
@@
To verify the effectiveness of the latent modulation, we compare the performance of different combinations of latent modulation. The results are shown in Table V. We can observe that the model using both scale modulation and shift modulation achieves the best performance, indicating that both types of modulation contribute effectively to enhancing arbitrary-scale SR performance.
V Conclusion
@@
In this paper, we propose the salient region-guided spacecraft image arbitrary-scale super-resolution network (SGSASR), which uses the feature information from the spacecraft core salient region to guide the generation of latent modulation and achieve arbitrary-scale SR. In SGSASR, we design a spacecraft core region recognition block (SCRRB) to identify the core salient region in spacecraft images by saliency detection. Additionally, we present an adaptive-weighted feature fusion enhancement mechanism (AFFEM), which flexibly fuses the spacecraft core region features with general image features using dynamic weight parameters. Extensive experiments on multiple datasets demonstrate the superior performance of SGSASR.
Acknowledgments
@@
The authors would like to thank Zhihui Li from China Aerodynamics Research and Development Center for providing the spacecraft ISAR data.

### v2

@@ 4.2 Comparison Results on Spacecraft Radar Image Dataset
Finally, we present visual results of spacecraft radar image SR using SGSASR at three non-integer scales ($\times$3.6, $\times$4.5, $\times$5.2) in Fig. 8. The input spacecraft image is downsampled by a factor of $\times$8 to generate the LR image.
Figure 8: Visual results of spacecraft radar image SR using SGSASR at non-integer scales. (a): LR image. (b): $\times$3.6. (c): $\times$4.5. (d): $\times$5.2.
Table 3: Quantitative comparisons with other fixed-scale SR methods on spacecraft optical image dataset. The best and the second best results are marked in bold and underline.
Methods $\times$4 $\times$8
$uparrow PSNR\uparrow SSIM\uparrow$
SSRN [9] 33.11 0.6967 31.99 0.6779
SSNet [35] 32.56 0.6607 31.30 0.6494
VapSR [20] 32.43 0.6876 30.64 0.6501
LKFN [22] 32.96 0.7011 31.73 0.6813
ELAN [25] 32.83 0.6948 31.88 0.6705
MAT [27] 32.52 0.6891 31.00 0.6550
SRConvNet [28] 32.95 0.6908 31.56 0.6594
SGSASR(Ours) 33.57 0.7132 32.09 0.6708
Table 4: Quantitative comparisons with other arbitrary-scale SR methods on spacecraft optical image dataset. The best and the second best results are marked in bold and underline.
@@
$uparrow PSNR\uparrow SSIM\uparrow$ $uparrow PSNR\uparrow SSIM\uparrow$
LIIF [11] 33.49 0.7073 32.63 0.6772 32.01 0.6634 31.10 0.6462 3723.64G
LM-LIIF [39] 33.44 0.7073 32.58 0.6788 31.97 0.6647 31.07 0.6484 486.64G
LM-LTE [39] 33.54 0.7074 32.68 0.6764 32.06 0.6613 31.14 0.6446 499.62G
SGSASR(Ours) 33.57 0.7132 32.72 0.6851 32.09 0.6708 31.17 0.6531 99.62G
Figure 9: Visual comparison for fixed-scale SR methods on spacecraft optical image dataset. (a): $\times 4$. (b): $\times 8$.
Figure 10: Visual comparison for arbitrary-scale SR methods on spacecraft optical image dataset. (a): $\times 4$. (b): $\times 6$. (c): $\times 8$. (d): $\times 12$.
4.3 Comparison Results on Spacecraft Optical Image Dataset
In this section, we compare SGSASR with the ten methods mentioned in Section 4.2 on the spacecraft optical image dataset. All methods are retrained on the spacecraft optical image dataset for a fair comparison. The quantitative comparison results with fixed-scale SR methods are presented in Table 3. The quantitative comparison results with arbitrary-scale SR methods are presented in Table 4. We can observe that SGSASR still outperforms fixed-scale SR methods in terms of PSNR performance on the spacecraft optical image dataset. Specifically, compared with SSRN, SGSASR achieves improvements of 0.46 dB and 0.10 dB at two SR scales ($\times$4, $\times$8), respectively. We can also observe that SGSASR consistently achieves superior PSNR performance at various SR scales on the spacecraft optical image dataset while maintaining lower computational costs compared with other arbitrary-scale SR methods.
In addition, we present visual comparison results for fixed-scale SR methods in Fig. 9 and for arbitrary-scale SR methods in Fig. 10. Overall, compared with other fixed-scale SR methods, SGSASR achieves comparable or even better restoration of spacecraft structural details on the spacecraft optical image dataset. In contrast to other arbitrary-scale SR methods, the spacecraft optical SR images generated by SGSASR not only more effectively restore the structure of the spacecraft, but also avoid introducing noise into the black background regions, which clearly demonstrates the superiority of SGSASR.
Finally, we present visual results of spacecraft optical image SR using SGSASR at three non-integer scales ($\times$3.6, $\times$4.5, $\times$5.2) in Fig. 11. The input spacecraft image is downsampled by a factor of $\times$8 to generate the LR image.
Figure 11: Visual results of spacecraft optical image SR using SGSASR at non-integer scales. (a): LR image. (b): $\times$3.6. (c): $\times$4.5. (d): $\times$5.2.
Table 5: Effect of Proposed Modules in SFEEM
@@ 4.3 Comparison Results on Spacecraft Optical Image Dataset
Figure 12: Visualization result of SCRRB. (a): The original spacecraft LR image. (b): The feature map extracted by SCRRB.
Table 6: Effect of Feature Aggregation Methods
@@
Table 7: Effect of Lantent Modulation in LMASSR
@@
4.4 Ablation Studies
4.4.1 The Effect of Proposed Modules in SFEEM
@@
To verify the effectiveness of the two main contributions proposed in this paper—–SCRRB and AFFEM, we conduct the incremental ablation experiments on the spacecraft ISAR image dataset. The results are shown in Table 5. The first row describes the Baseline model, which directly uses NAFNet [45] as the spacecraft image feature encoder. In the second row, we add the SCRRB, which yields a performance gain of 0.27 dB, since the SCRRB can better identify the core salient region in spacecraft images. In the third row, we further add the AFFEM, which yields an additional performance gain of 0.08 dB, since AFFEM can aggregate the spacecraft region features with general image features more flexibly.
4.4.2 The Visualization Result of SCRRB
To demonstrate the recognition ability of our SCRRB for the spacecraft core region, we provide the visualization of the feature map extracted by SCRRB in Fig. 12. We can observe that the response of the spacecraft body in the feature map is significantly stronger than that of the surrounding black background, indicating that SCRRB can effectively identify the spacecraft core salient region in the image.
4.4.3 The Effect of Feature Aggregation Methods
@@
To verify the effectiveness of our AFFEM, we compare it with other feature aggregation methods such as summation and concatenation. The results are shown in Table 6. We can observe that AFFEM achieves the best performance, demonstrating its effectiveness in adaptively aggregating the spacecraft region features with general image features.
4.4.4 The Effect of the Lantent Modulation in LMASSR
@@
To verify the effectiveness of latent modulation, we compare the performance of different combinations of latent modulation. The results are shown in Table 7. We can observe that the model using both scale modulation and shift modulation achieves the best performance, indicating that both types of modulation contribute effectively to enhancing arbitrary-scale SR performance.
5 Conclusion
@@
This paper introduces SGSASR, an efficient salient region-guided network for spacecraft image arbitrary-scale super-resolution. By leveraging the spacecraft core salient region to guide latent modulation, SGSASR effectively enhances SR performance. In SGSASR, we design a spacecraft core region recognition block (SCRRB) to identify the core salient region in spacecraft images by saliency detection. Additionally, we present an adaptive-weighted feature fusion enhancement mechanism (AFFEM), which flexibly fuses the spacecraft core region features with general image features using dynamic weight parameters. Extensive experiments on multiple datasets demonstrate that our SGSASR not only achieves state-of-the-art performance on spacecraft image arbitrary-scale SR while maintaining high computational efficiency, but also achieves competitive performance compared to fixed-scale SR methods.
This paper mainly studies typical spacecraft images with black space background, and has not yet considered spacecraft images with more complex backgrounds such as the Earth background. Therefore, in future work, we will develop arbitrary-scale super-resolution methods for spacecraft images with more challenging backgrounds.
CRediT authorship contribution statement
Jingfan Yang:Conceptualization, Data curation, Investigation, Methodology, Software, Writing – original draft. Hu Gao: Conceptualization, Formal analysis, Writing – review & editing. Ying Zhang: Investigation, Validation, Conceptualization. Depeng Dang: Funding acquisition, Supervision, Validation, Writing – review & editing.
Data availability
Spacecraft optical image data will be made available on request.
Acknowledgement
@@
This work was supported by the National Key Research and Development Program of China (Grant 2020YFC1523303). The authors would like to thank Zhihui Li from China Aerodynamics Research and Development Center for providing the spacecraft radar image dataset.